# The EU AI Act, Stakeholder Needs, and Explainable AI: Aligning Regulatory Compliance in a Clinical Decision Support System

### v1

@@ 3.1 The Cross-functional Team
The team behind this contribution consists of five individuals. Two individuals represent a company developing the CDSS. One of them has a PhD in computer science and has extensive experience in requirements engineering and empirical software engineering. The other is a PhD student who focuses on the human-centered implementation of XAI methods for AI-based CDSSs. Human-centered XAI methods put the human and their desiderata at the center of the implementation of XAI methods. The daily work of these individuals takes place in the context of an interdisciplinary exchange with the clinicians in a publicly funded research project. They organized the workshops (see Section 4) with usability engineers from the company, but are not part of the stakeholders that have been analyzed.
Two individuals represent a research institute; one with a background as a judge at an appeals court as well as working for multiple governmental investigations regarding the introduction of new laws, the other with a PhD in computer science with a focus on model-driven engineering and natural language processing. Together they have assisted the Government’s Office during negotiation and implementation of the AI Act and have shared their insights with the EU Commission. They also assist a notified body in defining the implications of the AI Act to existing and possible new business opportunities, e.g., when is a notified body needed, how to show fulfillment within a conformity assessment procedure and to which concrete products is the AI Act applicable.
The fifth member of the author team is a university professor for information systems and AI. The main responsibility in relation to this contribution is as an external auditor to assess the validity and scientific soundness of the contribution and deliver both a technical and socio-technical viewpoint.
@@ 3.2 The Clinical Decision Support System
This is where the CDSS comes into play. It supports clinicians by predicting key indicators that guide the transfer decision—such as remaining LOS, mortality probability, and readmission probability—using vital signs like heart rate and oxygen saturation from the past 24 hours. In particular, when the AI’s predictions conflict with a clinician’s intuition, the interpretability of the system is crucial for reconsidering and validating the final decision. To generate these predictions, a DNN is trained on time-series patient data spanning the years 2017 to 2024. Notably, the system does not suggest which patients to move, but only shows the predictions to help clinicians make decisions themselves.
@@ 3.3 Legal Analysis
The legal analysis consists of mapping the different articles of the AI Act to the system under investigation – which technical and organizational requirements does the system imply? Then, who will be responsible for the fulfillment of the requirements? The details of the technical requirements are not as important for the analysis as their implications for the different roles. For the legal analysis, e.g., it is more important to know who is responsible for which aspects of data management (input data, validation data, etc.), than how to assess the relevance or completeness of data. As analyses emerge we visualize the concepts and their relationships since text has the drawback of enforcing a linear representation (see e.g. Figure 2(a)). The primary outcome of the legal analysis is given in Section 4.
At the same time we make decisions on what is in and out of scope. The latter is necessary since the AI Act consists of 113 articles that define roles, responsibilities, technical requirements, fines and so on. That is complemented by 13 annexes that supply further detail to the articles. To motivate and help with interpretation of the articles and the annexes, there are 180 recitals. Knowing what to include in an analysis is therefore an analysis in itself and something that is made easier as more is known about the system under consideration. In the case of this contribution we settle on focusing primarily on the responsibilities of the roles of the provider and deployer, leaving out the content related to what happens if a system does not comply. We also keep the level of detail abstract in terms of technical requirements as going into detail would yield enough content for multiple academic contributions per article of the AI Act.
In parallel we conduct analyses of the AI Act in relation to the new machinery regulation as an assignment for the notified body (c) and a first analysis of the cyber-security requirements of the AI Act in relation to the regulation on medical devices and the Cyber Resilience Act (anonymized). These parallel investigations also suffice to better understand the AI Act in relations to the CDSS.
@@ 3.4 Requirements Engineering Activities
To understand the requirements of the clinicians that are going to use the CDSS, we conduct a workshop in which we collect the most important functional and quality requirements. The workshop is modeled after the Design Thinking method (Ferreira Martins et al., 2019), but much abbreviated. Participants were clinicians and ML researchers, with a requirements engineer moderating the workshop.
We engage in brainstorming sessions in which we map out the processes in the ICU and identify the most relevant user stories. These are then transformed into a user journey using user journey mapping (Milicic et al., 2012) to capture the requirements on the system implied by these processes. A scoping exercise then allows us to narrow down which aspects of the process can be supported by the system and thus define the scope of the CDSS.
To refine the requirements and the desiderata for the XAI methods in particular, we conduct several interviews with clinicians. We also perform contextual inquiry (Bednar and Welch, 2009) where we follow a clinician on their rounds in the ICU to better understand how the process we elicit works in practice and which factors they consider before deciding which patient to move out of the ICU.
@@ 3.5 Stakeholder Mapping and XAI Analysis
The next step of analysis is to identify the relevant stakeholder groups. Thereby, we have discussions involving the AI Act and XAI experts, following the stakeholder mapping approach outlined by Langer et al. (2021). Stakeholders identified during the Design Thinking Workshop are mapped, and we summarize their specific desiderata and expand upon them in regular project meetings and exchanges. For the choice of XAI methods, the categorization by Speith (2022) is used. The Ph.D. student analyzes the problem and selects common methods that align with the problem types and relevant requirements of the CDSS, as identified in the requirements engineering workshops.
@@ 3.6 The Merging of the two Approaches
The legal and the XAI analysis are made in parallel so that they benefit from insights from the other approach. This takes place over three different phases. Phase one was a one-day sprint (March 2024) where the AI Act is introduced to the company and a subsequent workshop starts to map out the high-level applicability towards the CDSS. Phase two occurs six months later (September 2024) to share new insights in a second workshop. The third phase runs from December 2024 to January 2025 and is conducted as six weekly meetings with the ambition to document the combined analysis and the reasoning. Between the meetings of the third phase, each individual contributes to the shared document and the meetings discuss the combined effort and changes are made as new insights were gained. The outcome in this contribution represents the final and polished analyses.
@@ 3.7 Validity Strategies
Following Lincoln and Guba (1985), we structure the validity strategies along Credibility (establishing an analysis true to the data and context); Transferability (ensuring that the findings were applicable beyond the context); Dependability (enabling repetition of our contribution); and Confirmability (limiting the impact of our own incentives and bias).
As seen in Section 3.6, we apply a methodology that allows for triangulation of methodologies applied separately by the individuals of the team, over time (Credibility and Confirmability). We also show how our analysis goes beyond a specific AI system in sections 6 and 7 (Transferability). The ambition is that it is possible to replicate the study as we describe the team behind the contribution and its incentives, as well as the applied methodologies for collecting and analyzing data (Dependability and Confirmability).
@@ 4 Applicability of the AI Act to the CDSS
A key concept here is infer which can be done by machine-learning approaches but also through knowledge- or logic-based technologies. However, “systems that are based on the rules defined solely by natural persons to automatically execute operations” should not be seen as AI systems (recital 12). Our assessment is that the AI system under investigation is within scope of the definition of AI system as it, from received input, generates a prediction that, if the clinician approves, will influence where and by who the patient will be treated.
@@
Conclusion: The AI system needs to be CE-marked through an internal conformity assessment procedure in accordance with the AI Act. This is the obligation of the provider. In practice, this means that the provider needs to fulfill several obligations, and described in Section 2.1. This includes fulfilling technical criteria (Art. 9-15) and having a quality management system (Art. 17).
That is mirrored by the obligations of the deployer (Art. 26), also described in Section 2.1. Since the system will be deployed by a legal person governed by public law (or by private entities providing public services), there is another obligation: a Fundamental Rights Impact Assessment shall be conducted by the deployer (Art. 27). How to do this exactly will be further detailed in a harmonized standard.
@@ 5 Choosing the right XAI Method
Choosing the right XAI method is important to ensure that the stakeholder needs are satisfied (Schoonderwoerd et al., 2021). Therefore, we analyze relevant stakeholders and elaborate their desiderata (see Section 5.1). This information helps us make well-founded choices about the XAI method in the implementation phase of the CDSS (see Section 5.2).
@@ 5.1 Our Stakeholders and their Desiderata
The first group of stakeholders are researchers and engineers specializing in ML, who are categorized as developers and identified as providers under the provisions of the AI Act. Their primary requirement for the AI system is the achievement of high performance and reliability. In order to fulfill this requirement, they concentrate on quality measures such as accuracy and robustness, which are of the highest importance for the practical implementation and success of the model in real-world applications. This will increase end-users’ acceptance of our AI system which is a key factor for hospital management.
As the deployer of the AI system, the hospital management is interested in acceptance of the AI system by the clinicians. It needs to be a reliable source of support in high-stakes decision-making processes. Furthermore, the AI system must comply with a range of regulations, such as the AI Act, to ensure its appropriate in the hospital.
Clinicians are the primary users of the AI system in CDSS, as they are the only individuals who directly interact with the AI outputs and ultimately make the final decisions. Most of these clinicians are novice users; as defined in Section 2.2, the AI system must be accessible and user-friendly. This translates into the desideratum interpretability, which involves providing comprehensible, plausible, and correct explanations for the AI system. This interpretability allows clinicians to question AI predictions rather than simply accepting them, thereby fostering trust and appropriate reliance (Schemmer et al., 2023). Clinicians should trust the AI system when it is correct and appropriately rely on their judgment when the AI system is incorrect. This is most important when the AI system makes a false prediction that negatively impacts the patient, e.g., the AI system falsely predicts a short remaining LOS for the patients, although the patient is very sick and will have to stay longer. In this situation, the clinicians must rely on their medical expertise.
Patients, although not directly interacting with the AI system, are significantly impacted by the decisions made by clinicians using the AI system and thus an “affected party” according to the definition in Section 2.2. Patients have a strong need for correct and fair decision-making. Fairness in AI-driven decisions is a critical requirement, as it directly affects patient outcomes and trust in the healthcare system.
The regulator corresponds in our case to the competent authority, which can be mapped to different AI Act stakeholder roles, e.g.: AI Office or AI Board.
@@ 5.2 Our XAI Method Choices
After we clarified the relevant stakeholders and their desiderata, we now describe the concrete XAI method choices we made for our CDSS. We made choices in the categories applicability, scope, and stage, as proposed by Speith (2022). Categories like input data and problem type are pre-determined by the problem the application addresses.
@@
Stage: While DNNs provide great performance by learning complex relationships in data, they are not intrinsically interpretable and lack transparency (Barnes and Hutson, 2024). Therefore, we employ post-hoc XAI methods to provide insights into this black-box model.
@@
Finally, the different choices in the categories allow us to choose between XAI methods such as Shap (Lundberg and Lee, 2017), LIME (Ribeiro et al., 2016), Integrated Gradient (Sundararajan et al., 2017), or Saliency Maps (Simonyan et al., 2014). Functionally-grounded evaluation methods (Doshi-Velez and Kim, 2017) that measure explanation properties such as complexity, faithfulness, or sensitivity help us to select the best XAI method (Chen et al., 2024).
For example, Figure 1 shows a concrete heatmap explanation highlighting the importance of features for all time series steps. Here, the clinician gets an overview of the five most important features over the last 15 hours of the patient in the ICU. The color gradient indicates if the specific feature value at the specific time stamp has a negative (green) or positive (red) effect on the predicted remaining LOS. A short LOS indicates a more positive state of health then a long LOS. This explanation of the LOS prediction assists the clinician in appropriately deciding whether the patient can be moved to another ward.
@@ 6 Contrasting XAI with the AI Act
The AI Act refers to stakeholder obligations, while XAI describes stakeholder desiderata. Mapping these two perspectives is not a straightforward one-to-one mapping. Rather than exhaustively detailing the inter-relationships, we explored a thin slice of their intersection to show how XAI can fulfill the AI Act’s legal requirements. This means we are explicitly expecting further research to complement and refute our findings (see KI1-KI3), both in scope and in conclusions if XAI is chosen as a relevant way to comply with the requirements of the AI Act. A schematic view of the analysis is found in diagrams of Figure 2.
@@ 6.1 Roles, Responsibilities and Needs
As laid out in Section 2.2, the XAI community is concerned with stakeholder desiderata (Langer et al., 2021) while the AI Act concerns stakeholder responsibilities (see Section 2.1). While there is an overlap in the main stakeholder roles introduced by XAI and the AI Act, there are also differences. Where the XAI community is satisfied by the role of regulator (Langer et al., 2021), the AI Act has multiple roles. There is the national authority assigned to conduct market surveillance of AI systems, the notified body that inspects to what extent an AI system and the developer’s organization conform to the requirements and the EU Commission who all have a say in how the AI Act is to be implemented and enforced. Even a court could serve as a regulator if the system developer and the national authority have different opinions. In the case of our CDSS, the system developer can even take the XAI role of regulator. This can seem counter-intuitive, but it is one way of interpreting the conformity procedure based on self-assessment.
On the other side, the XAI community pays explicit attention to the end user and the affected parties (Langer et al., 2021), while the AI Act trusts the provider and the deployer to take care of those perspectives. As such the stakeholder roles can be seen as lacking a common ground, or to complement each other. That leaves ample room for the XAI community to define their stakeholder roles in relation to the responsibilities enforced by the AI Act.
Another observation is that while Langer et al. (2021) treat individuals “as representative members of specified stakeholder classes”, our analysis mandates us to distinguish between affected parties as concrete individuals and as an abstract collective due to the responsibilities of the deployer. Art. 26 in the AI Act defines requirements on the usage of the AI system on an individual level at run-time, while Art. 27 poses requirements on performing risk assessments of the AI system in relation to its intended usage pre-launch in respect to the affected parties in general. This is visualized in Figure 2(a) as two different responsibilities (arrows) between the deployer and the affected parties.

### v2

@@ 3.1 The Cross-functional Team
The team behind this contribution consists of five individuals. Two individuals represent a company developing the CDSS. One of them has a PhD in computer science and has extensive experience in requirements engineering and empirical software engineering. The other is a PhD student who focuses on the human-centered implementation of XAI methods for AI-based CDSSs. Human-centered XAI methods put the human and their desiderata at the center of the implementation of XAI methods. The daily work of these individuals takes place in the context of an interdisciplinary exchange with the clinicians in a publicly funded research project. They organized the workshops (see Section˜4) with usability engineers from the company, but they are not part of the stakeholders that have been analyzed.
Two individuals represent a research institute; one with a background as a judge at an appeals court, as well as working for multiple governmental investigations regarding the introduction of new laws, and the other with a PhD in computer science with a focus on model-driven engineering and natural language processing. Together, they have assisted the Government’s Office during the negotiation and implementation of the AI Act and have shared their insights with the EU Commission. They also assist a notified body in defining the implications of the AI Act to existing and possible new business opportunities, e.g., when a notified body is needed, how to show fulfillment within a conformity assessment procedure, and to which concrete products the AI Act is applicable.
The fifth member of the author team is a university professor for information systems and AI. The primary responsibility in relation to this contribution is to serve as an external auditor, assessing the validity and scientific soundness of the contribution and delivering both a technical and socio-technical viewpoint.
@@ 3.2 The Clinical Decision Support System
This is where the CDSS comes into play. It supports clinicians by predicting key indicators that guide the transfer decision—such as remaining LOS, mortality probability, and readmission probability—using vital signs like heart rate and oxygen saturation from the past 24 hours. In particular, when the AI’s predictions conflict with a clinician’s intuition, the interpretability of the system is crucial for reconsidering and validating the final decision. To generate these predictions, a DNN is trained on time-series patient data spanning the years 2017 to 2024. Notably, the system does not suggest which patients to move, but rather shows the predictions to help clinicians make their own decisions.
@@ 3.3 Legal Analysis
The legal analysis consists of mapping the different articles of the AI Act to the system under investigation—which technical and organizational requirements does the system imply? Then, who will be responsible for the fulfillment of the requirements? The details of the technical requirements are not as important for the analysis as their implications for the different roles. For the legal analysis, e.g., it is more important to know who is responsible for which aspects of data management (input data, validation data, etc.), than how to assess the relevance or completeness of data. As analyses emerge, we visualize the concepts and their relationships since text has the drawback of enforcing a linear representation (see e.g. Figure˜2(a)). The primary outcome of the legal analysis is given in Section˜4.
At the same time, we make decisions on what is in and out of scope. The latter is necessary since the AI Act consists of 113 articles that define roles, responsibilities, technical requirements, fines, and so on. That is complemented by 13 annexes that supply further detail to the articles. To motivate and help with the interpretation of the articles and the annexes, there are 180 recitals. Knowing what to include in an analysis is therefore an analysis in itself and something that is made easier as more is known about the system under consideration. In the case of this contribution, we settle on focusing primarily on the responsibilities of the roles of the provider and deployer, leaving out the content related to what happens if a system does not comply. We also keep the level of detail abstract in terms of technical requirements as going into detail would yield enough content for multiple academic contributions per article of the AI Act.
In parallel, we conduct analyses of the AI Act in relation to the new machinery regulation as an assignment for the notified body (c) and a first analysis of the cybersecurity requirements of the AI Act in relation to the regulation on medical devices and the Cyber Resilience Act (anonymized). These parallel investigations also suffice to better understand the AI Act in relation to the CDSS.
@@ 3.4 Requirements Engineering Activities
To understand the requirements of the clinicians who are going to use the CDSS, we conduct a workshop in which we collect the most important functional and quality requirements. The workshop is modeled after the Design Thinking method (ferreira2019design), but much abbreviated. Participants were clinicians and ML researchers, with a requirements engineer moderating the workshop.
We engage in brainstorming sessions in which we map out the processes in the ICU and identify the most relevant user stories. These are then transformed into a user journey using user journey mapping (milicic2012towards) to capture the requirements on the system implied by these processes. A scoping exercise then allows us to narrow down which aspects of the process can be supported by the system and thus define the scope of the CDSS.
To refine the requirements and the desiderata for the XAI methods in particular, we conduct several interviews with clinicians. We also perform contextual inquiry (bednar2009contextual) where we follow a clinician on their rounds in the ICU to better understand how the process we elicit works in practice and which factors they consider before deciding which patient to move out of the ICU.
@@ 3.5 Stakeholder Mapping and XAI Analysis
The next step of analysis is to identify the relevant stakeholder groups. Thereby, we have discussions involving the AI Act and XAI experts, following the stakeholder mapping approach outlined by langer_what_2021. Stakeholders identified during the Design Thinking Workshop are mapped, and we summarize their specific desiderata and expand upon them in regular project meetings and exchanges. For the choice of XAI methods, the categorization by speith_review_2022 is used. The Ph.D. student analyzes the problem and selects common methods that align with the problem types and relevant requirements of the CDSS, as identified in the requirements engineering workshops.
@@ 3.6 The Merging of the two Approaches
The legal and the XAI analysis are made in parallel so that they benefit from insights from the other approach. This takes place over three different phases. Phase one was a one-day sprint (March 2024) where the AI Act is introduced to the company and a subsequent workshop starts to map out the high-level applicability towards the CDSS. Phase two occurs six months later (September 2024) to share new insights in a second workshop. The third phase runs from December 2024 to January 2025 and is conducted as six weekly meetings with the ambition to document the combined analysis and the reasoning. Between the meetings of the third phase, each individual contributes to the shared document, and the meetings discuss the combined effort, and changes are made as new insights are gained. The outcome in this contribution represents the final and polished analyses.
@@ 3.7 Validity Strategies
Following LincolnGuba1985, we structure the validity strategies along Credibility (establishing an analysis true to the data and context); Transferability (ensuring that the findings were applicable beyond the context); Dependability (enabling repetition of our contribution); and Confirmability (limiting the impact of our own incentives and bias).
As seen in Section˜3.6, we apply a methodology that allows for triangulation of methodologies applied separately by the individuals of the team, over time (Credibility and Confirmability). We also show how our analysis goes beyond a specific AI system in sections 6 and 7 (Transferability). The ambition is that it is possible to replicate the study as we describe the team behind the contribution and its incentives, as well as the applied methodologies for collecting and analyzing data (Dependability and Confirmability).
@@ 4 Applicability of the AI Act to the CDSS
A key concept here is infer, which can be done by machine-learning approaches but also through knowledge- or logic-based technologies. However, “systems that are based on the rules defined solely by natural persons to automatically execute operations” should not be seen as AI systems (recital 12). Our assessment is that the AI system under investigation is within the scope of the definition of an AI system as it, from received input, generates a prediction that, if the clinician approves, will influence where and by whom the patient will be treated.
@@
Conclusion: The AI system needs to be CE-marked through an internal conformity assessment procedure in accordance with the AI Act. This is the obligation of the provider. In practice, this means that the provider needs to fulfill several obligations, as described in Section˜2.1. This includes fulfilling technical criteria (Art. 9-15) and having a quality management system (Art. 17).
That is mirrored by the obligations of the deployer (Art. 26), also described in Section˜2.1. Since the system will be deployed by a legal person governed by public law (or by private entities providing public services), there is another obligation: a Fundamental Rights Impact Assessment shall be conducted by the deployer (Art. 27). How to do this exactly will be further detailed in a harmonized standard.
@@ 5 Choosing the right XAI Method
Choosing the right XAI method is important to ensure that the stakeholder needs are satisfied (schoonderwoerd_human-centered_2021). Therefore, we analyze relevant stakeholders and elaborate their desiderata (see Section˜5.1). This information helps us make well-founded choices about the XAI method in the implementation phase of the CDSS (see Section˜5.2).
@@ 5.1 Our Stakeholders and their Desiderata
The first group of stakeholders are researchers and engineers specializing in ML, who are categorized as developers and identified as providers under the provisions of the AI Act. Their primary requirement for the AI system is the achievement of high performance and reliability. In order to fulfill this requirement, they concentrate on quality measures such as accuracy and robustness, which are of the highest importance for the practical implementation and success of the model in real-world applications. This will increase end-users’ acceptance of our AI system, which is a key factor for hospital management.
As the deployer of the AI system, the hospital management is interested in the acceptance of the AI system by the clinicians. It needs to be a reliable source of support in high-stakes decision-making processes. Furthermore, the AI system must comply with a range of regulations, such as the AI Act, to ensure its appropriateness in the hospital.
Clinicians are the primary users of the AI system in CDSS, as they are the only individuals who directly interact with the AI outputs and ultimately make the final decisions. Most of these clinicians are novice users; as defined in Section˜2.2, the AI system must be accessible and user-friendly. This translates into the desideratum interpretability, which involves providing comprehensible, plausible, and correct explanations for the AI system. This interpretability allows clinicians to question AI predictions rather than simply accepting them, thereby fostering trust and appropriate reliance (schemmer_appropriate_2023). Clinicians should trust the AI system when it is correct and appropriately rely on their judgment when the AI system is incorrect. This is most important when the AI system makes a false prediction that negatively impacts the patient, e.g., the AI system falsely predicts a short remaining LOS for the patients, although the patient is very sick and will have to stay longer. In this situation, the clinicians must rely on their medical expertise.
Patients, although not directly interacting with the AI system, are significantly impacted by the decisions made by clinicians using the AI system and thus an “affected party” according to the definition in Section˜2.2. Patients have a strong need for correct and fair decision-making. Fairness in AI-driven decisions is a critical requirement, as it directly affects patient outcomes and trust in the healthcare system.
The regulator corresponds in our case to the competent authority, which can be mapped to different AI Act stakeholder roles, e.g., AI Office or AI Board.
@@ 5.2 Our XAI Method Choices
After we clarified the relevant stakeholders and their desiderata, we now describe the concrete XAI method choices we made for our CDSS. We made choices in the categories applicability, scope, and stage, as proposed by speith_review_2022. Categories like input data and problem type are pre-determined by the problem the application addresses.
@@
Stage: While DNNs provide great performance by learning complex relationships in data, they are not intrinsically interpretable and lack transparency (barnes_evaluating_2024). Therefore, we employ post-hoc XAI methods to provide insights into this black-box model.
@@
Finally, the different choices in the categories allow us to choose between XAI methods such as Shap (lundberg_unified_2017), LIME (ribeiro_why_2016), Integrated Gradient (sundararajan_axiomatic_2017), or Saliency Maps (simonyan_deep_2014). Functionally-grounded evaluation methods (doshi-velez_towards_2017) that measure explanation properties such as complexity, faithfulness, or sensitivity help us to select the best XAI method (chenWhatMakesGood2024).
For example, Figure˜1 shows a concrete heatmap explanation highlighting the importance of features for all time series steps. Here, the clinician gets an overview of the five most important features over the last 15 hours of the patient in the ICU. The color gradient indicates if the specific feature value at the specific time stamp has a negative (green) or positive (red) effect on the predicted remaining LOS. A short LOS indicates a more positive state of health then a long LOS. This explanation of the LOS prediction assists the clinician in appropriately deciding whether the patient can be moved to another ward.
@@ 6 Contrasting XAI with the AI Act
The AI Act refers to stakeholder obligations, while XAI describes stakeholder desiderata. Mapping these two perspectives is not a straightforward one-to-one mapping. Rather than exhaustively detailing the inter-relationships, we explored a thin slice of their intersection to show how XAI can fulfill the AI Act’s legal requirements. This means we are explicitly expecting further research to complement and refute our findings (see KI1-KI3), both in scope and in conclusions, if XAI is chosen as a relevant way to comply with the requirements of the AI Act. A schematic view of the analysis is found in diagrams of Figure˜2.
@@ 6.1 Roles, Responsibilities and Needs
As laid out in Section˜2.2, the XAI community is concerned with stakeholder desiderata (langer_what_2021) while the AI Act concerns stakeholder responsibilities (see Section˜2.1). While there is an overlap in the main stakeholder roles introduced by XAI and the AI Act, there are also differences. Where the XAI community is satisfied by the role of regulator (langer_what_2021), the AI Act has multiple roles. There is the national authority assigned to conduct market surveillance of AI systems, the notified body that inspects to what extent an AI system and the developer’s organization conform to the requirements and the EU Commission who all have a say in how the AI Act is to be implemented and enforced. Even a court could serve as a regulator if the system developer and the national authority have different opinions. In the case of our CDSS, the system developer can even take the XAI role of regulator. This can seem counterintuitive, but it is one way of interpreting the conformity procedure based on self-assessment.
On the other side, the XAI community pays explicit attention to the end user and the affected parties (langer_what_2021), while the AI Act trusts the provider and the deployer to take care of those perspectives. As such, the stakeholder roles can be seen as lacking a common ground, or as complementing each other. That leaves ample room for the XAI community to define their stakeholder roles in relation to the responsibilities enforced by the AI Act.
Another observation is that while langer_what_2021 treat individuals “as representative members of specified stakeholder classes”, our analysis mandates us to distinguish between affected parties as concrete individuals and as an abstract collective due to the responsibilities of the deployer. Art. 26 in the AI Act defines requirements on the usage of the AI system on an individual level at run-time, while Art. 27 poses requirements on performing risk assessments of the AI system in relation to its intended usage pre-launch in respect to the affected parties in general. This is visualized in Figure˜2(a) as two different responsibilities (arrows) between the deployer and the affected parties.